OpenStudy (goformit100):

If 1+9 = 11 then 2 +10 = ?

OpenStudy (goformit100):

@Nanoman @nader1

OpenStudy (goformit100):

@Eyad

OpenStudy (goformit100):

@Zale101

OpenStudy (anonymous):

1+9=10

OpenStudy (anonymous):

2+10=12

OpenStudy (anonymous):

Or are my calculations wrong?

OpenStudy (zale101):

this question is REALLY tough.

OpenStudy (zale101):



OpenStudy (zale101):

2+10=13

OpenStudy (anonymous):



OpenStudy (anonymous):

Is this calculus or something because I always thought 1 + 1 =2